the permiter of a rectangular playing field is 402 yards. The length of the field is 7 yards less than triple the width. what are the dimensions of the playing field
Answer by addingup(3677)

2Length + 2Width = Perimeter
Length = 3W - 7
2(3W - 7) + 2W = 402
6W - 14 + 2W = 402
8W = 416
W = 52
and the Length is 3W - 7 = 3(52) - 7 = 149
------------------------
Check:
2Length + 2 Width = Perimeter
2(149) + 2(52) = 402
298 + 104 = 402 Correct